OpenStudy (anonymous):

What is the value of y when x = 4? 2x - 3y = 20 -9 -4 4 8

OpenStudy (texaschic101):

2x - 3y = 20 when x = 4 just sub in 4 for x and solve for y 2(4) - 3y = 20 8 - 3y = 20 == subtract 8 from both sides -3y = 20 - 8 -3y = 12 == divide both sides by -3 y = 12/-3 y = -4